OpenStudy (anonymous):

what is the axis of symmetry for the equation x= x^2+8x+3 ?

OpenStudy (amistre64):

x= or y=?

OpenStudy (anonymous):

sorry i meant f(x)=

OpenStudy (amistre64):

ah ....

OpenStudy (amistre64):

do you know how to find the vertex? or to put this into f(x)=a(x-h)^2+k format?

OpenStudy (anonymous):

not really..

OpenStudy (amistre64):

hmm, another way of looking at it would be if you know how to find the zeros, or roots, of the equation; the line of symmetry would be the midpoint between the roots .... provided it crosses the x axis to begin with

OpenStudy (amistre64):

formula-wise we have: given: ax^2 + bx + c the line of symmetry can be determined as: x = -b/2a

OpenStudy (anonymous):

so im going to put -8/ 2(1) ?

OpenStudy (amistre64):

thats perfect

OpenStudy (anonymous):

oh yay! so i get x=-4?

OpenStudy (amistre64):

yes, x=-4 will be the axis of symmetry, good job

OpenStudy (anonymous):

so how do i find the vertex?

OpenStudy (amistre64):

the vertex actully sits right on the axis of symmetry; so we know x=-4 is the axis; use x=-4 in the equation itself to find the y value that defines the point for the vertex

OpenStudy (amistre64):

otherwise we would have to go the route of completing the square .... which isnt difficult, but is rather involved

OpenStudy (anonymous):

wait which equation?

OpenStudy (amistre64):

x^2+8x+3 the one youre asking about is this one correct?

OpenStudy (anonymous):

yes! so i'm going to put -4 in each x?

OpenStudy (amistre64):

correct, since we know the x part of the point has to be -4

OpenStudy (anonymous):

so one of the points are going to be -4? the x one? & when i plug the -4 into the x's its going to give me the y? am i right or way off?

OpenStudy (amistre64):

that is correct.

OpenStudy (anonymous):

i got -13 when i plugged in the -4 to all the x's!

OpenStudy (anonymous):

so my vertex is (-4, -13) ?

OpenStudy (amistre64):

lets see what i get to verify 16-32+3 = -16+3 = -13 yes

OpenStudy (amistre64):

correct (-4,-13) is the vertex of this parabola then

OpenStudy (anonymous):

okay and since its both negative, the parabola is going to be like...

OpenStudy (anonymous):



OpenStudy (anonymous):

so minimum!

OpenStudy (amistre64):

not quite, the vertex only tells us where the parabola bends at; the direction that it open up towards is determined by looking at the first term of the eqaution; the "ax^2" part

OpenStudy (amistre64):

ax^2 is positive so it opens up -ax^2 is negtaive so it opens down

OpenStudy (anonymous):

so its going to be going the other wayy?

OpenStudy (amistre64):

given your equation; it starts with: x^2 so it will open the same way as x^2 does: like a U

OpenStudy (anonymous):

so thats a maximum?

OpenStudy (amistre64):

if your question is whether the vertex is the highest (max) point or the lowest (min) point on the parabola; i would go with lowest since all the other points are moving up and higher than it

OpenStudy (anonymous):

it says state whether the vertex is max or min..

OpenStudy (amistre64):